# “6 choose 4”: A framework to understand and facilitate discussion of strategies for overall survival safety monitoring

### v1

@@ 3 Case study
Rodriguez et al. (2024) considered controlling type I error in the strong sense. For Strategies 2A and 2B, the family wise error rates (i.e., probability of meeting at least one threshold under $H_{0}$) are 0.049 and 0.12, respectively (not shown in Table 1). Therefore, if there is a desire to control these rates at a specified level, say 0.025 or 0.10, then one can expect probabilities less than 60% and 78.3% of meeting all three thresholds under $H_{1}$—levels of risk that sponsors might be unwilling to take. In the context of safety monitoring where interest is in flagging a treatment that causes OS detriment, we think the probability of consistently being deemed safe under $H_{0}$—or more precisely, its complement, the probability of being flagged as unsafe at least once if the drug is harmful—is a more meaningful quantity than FWER.
@@ 4 Discussion
We described five special cases under the proposed framework: Fleming et al. (2024), Rodriguez et al. (2024), Shan (2023), an approach that uses only discrete thresholds for decision making, and another approach that is based off of previous FDA monitoring guidelines for evaluating cardiovascular risk in T2D. We illustrated use of the five approaches by applying them to the POLARIX trial (Section 3). While the main objective of our paper is to increase understanding of each approach, thereby facilitating planning by sponsors of future clinical trials and improving communication between stakeholders, we leave here with some final comments and recommendations regarding their use.
@@
Of the five approaches mentioned in this paper, we think Fleming et al. (2024) is the most appropriate for practical use due to its balance in priority over time—which shifts from a focus on ensuring that potentially effective treatments will continue to be studied when there are few OS events, to a focus on ruling out OS detriment when more OS events have accumulated. This approach also has the advantage of making clear how decision making thresholds can be re-calculated at any time given the observed number of OS events.
@@
Approach #5 based on FDA’s guidance in T2D can be used in tandem with Fleming et al. (2024) to better evaluate benefit-risk. Together, when deaths are still slowly accruing, the two approaches address the questions, “What OS detriment can we rule out with a high level of certainty?” (approach #5) and “With what certainty can we rule out a particular level of OS detriment?” (Fleming et al., 2024).
As for the remaining three approaches, users should use Rodriguez et al. (2024) with caution, as overemphasizing control of type I error early risks stopping the investigation of effective therapies prematurely. Users can choose to relax the marginal type I error rates, but more guidance is needed on how to do so. Shan (2023) cannot be used for earlier interim analyses whose timing is driven by an intermediate/surrogate endpoint, but it can be used for later interim analyses whose timings are driven by OS. Finally, use of discrete thresholds can be appropriate provided that the anticipated risks involved are understood and acceptable.
Nevertheless, for all three of these approaches, the actual type I and type II error rates from using the decision making thresholds will depend on the observed number of OS events. Therefore, sponsors should have a plan in place on how they will make adjustments in case the observed number of OS events at an analysis time deviates from the number of OS events anticipated at the planning stage. Alternatively, sponsors could re-frame their approach under Fleming et al. (2024) with similar operating characteristics. Measures of overall rates—such as the probability of a detrimental drug being flagged at least once for safety concern, or the probability that an effective drug is consistently deemed safe (Section 3)—can help to inform the choice of one approach over another, as well as to update the parameter values during trial conduct.
@@
Regardless of the approach taken, the low mortality rate in indolent diseases makes rigorous OS monitoring challenging. Intercurrent events such as subsequent therapy may further increase uncertainty by confounding the treatment effect (Korn et al., 2011). Therefore, multiple iterations of calculations and discussions are likely needed in practice for stakeholders to achieve alignment. We agree with Fleming et al. (2024) that any set of proposed thresholds should be seen as non-binding “informative guidelines” as opposed to black-and-white “rules” for continuing or stopping a trial. We find the list of additional analyses suggested by Rodriguez et al. (2024) to increase understanding of immature OS data helpful: analysis of biologically plausible subgroups, simulation and supplementary analyses to investigate the robustness of results under different assumptions, evaluation of other endpoints, results from other trials, use of real-world data, etc.. Data monitoring committees may also find measures of stability (or graphical summaries such as the predicted interval plot) useful for understanding potential effect size estimates and associated precision with trial continuation (Betensky, 2015; Li et al., 2009).
Several questions remain to be addressed for the practical implementation of OS safety monitoring. How should $\theta_{0}$ be defined, potentially taking into consideration the treatment effect that will be observed for other endpoints? (For example, a greater tolerance for OS risk might be more acceptable when a robust clinical benefit is ascertained on an intermediate/surrogate endpoint.) Likewise, how should $\theta_{1}$ be defined, given potential confounding by intercurrent events? How frequently should OS be evaluated for safety, i.e., how should we define the “analysis times” in the proposed framework? In what contexts should OS safety be formally tested and how should it be done relative to other formally tested endpoints—which may or may not include OS for efficacy?
@@
This paper proposes a supportive decision making framework based on a single criteria (i.e., $\theta^{*}$). However, it is worth noting that Rodriguez et al. (2024) and Shan (2023) considered introducing a ‘grey zone’, which, if a trial were to fall into, would prompt further OS follow-up. Doing so could have the effect of ruling out substantial OS detriment while limiting rejection of treatments with marginal OS benefit. While we did not consider such an extension here for simplicity of comparison and presentation, we do find this idea of practical interest worth pursuing in future discussions and research.

### v2

@@ 3 Case study
For Strategy 3A, the family wise error rate (i.e., probability of being deemed safe at least once under $H_{0}$) is 0.053 (not shown in Table 1). If there is a desire to control this rate at a lower level, say 0.05 or 0.025, then one can expect a less than 58% probability of consistently being deemed safe under $H_{1}$—a level of risk that sponsors might be unwilling to take. In the context of safety monitoring where interest is in flagging a treatment that causes OS detriment, we think the probability of consistently being deemed safe under $H_{0}$—or more precisely, its complement, the probability of being flagged as unsafe at least once if the drug is harmful—is a more meaningful quantity than the FWER.
@@ 4 Discussion
We described five special cases under the proposed framework: Fleming et al. (2024), Rodriguez et al. (2024), an approach that uses standard confidence intervals to continually rule out harm, another approach that uses discrete thresholds for decision making, and a fifth approach that is based off of previous FDA monitoring guidelines for evaluating cardiovascular risk in T2D. We illustrated use of the five approaches by applying them to the POLARIX trial (Section 3). While the main objective of our paper is to increase understanding of each approach, thereby facilitating planning by sponsors of future clinical trials and improving communication between stakeholders, we leave here with some final comments and recommendations regarding their use.
@@
Of the five approaches mentioned in this paper, we think Fleming et al.’s is the most appropriate for general use due to its balance in priority over time—which shifts from a focus on ensuring that potentially effective treatments will continue to be studied when there are few OS events, to a focus on ruling out OS detriment when more OS events have accumulated. This approach also has the advantage of making clear how decision making thresholds can be calculated at any time given the observed number of OS events.
@@
The fifth approach based on FDA’s guidance in T2D can be used in tandem with Fleming et al. to better evaluate benefit-risk. Together, when deaths are still slowly accruing, the two approaches address the questions, “What OS detriment can we rule out with a high level of certainty?” and “With what certainty can we rule out a particular level of OS detriment?”.
Rodriguez et al. can be used to identify the goal of a clinical trial, the point at which OS data can be considered sufficiently mature to facilitate evaluation of OS as a safety endpoint. However, it provides no guidance on how to evaluate OS at other time points, including potential interim analyses driven by the primary endpoint. Likewise, 95% CIs provide little utility for ruling out harm early on in a trial.
More relaxed CIs (90%, 80%, etc.) or discrete thresholds may be appropriate provided that the risks involved are understood and acceptable. In fact, one could apply Fleming et al. as a first approximation, then find an approach using standard CIs or discrete thresholds that has similar operating characteristics. Such an approach could effectively balance priorities over time and be easy to implement during a trial. Sponsors taking this approach would do well to pre-specify when and how to adjust the $\alpha$-level or discrete threshold in case the number of events deviates significantly from expectations during trial conduct (see Figure 1 for example).
Figure 1: Fleming et al. (2024) and an approximately equivalent discrete threshold strategy applied to POLARIX at the first planned interim analysis with 89 expected OS events. Solid curve is the hazard ratio (HR) threshold for decision making based on Fleming et al. (2024) with $\theta_{0}=1.3$, $\theta_{1}=0.80$, $1-\beta=0.9$, and varying potential number of observed deaths $d$. Dashed curve indicates discrete thresholds that approximate Fleming et al. (2024) by maintaining $1-\beta\geq 0.9$ while tolerating higher levels of $\alpha$.
@@
Regardless of the approach taken, the low mortality rate in indolent diseases makes rigorous OS monitoring challenging. We agree with Fleming et al. (2024) that any set of proposed thresholds should be seen as non-binding “informative guidelines” as opposed to black-and-white “rules” for continuing or stopping a trial. At a minimum, the need to monitor OS for safety should encourage stakeholders to contemplate the level of OS data that is likely to be accessible during the course of a clinical trial. Meanwhile, the additional analyses suggested by Rodriguez et al. (2024) may be helpful for gaining a better understanding of immature OS data: analysis of biologically plausible subgroups, simulation and supplementary analyses to investigate the robustness of results under different assumptions, evaluation of other endpoints, results from other trials, use of real-world data, etc.. IDMCs may also find measures of stability or graphical summaries such as the predicted interval plot useful for understanding potential effect size estimates and associated precision with trial continuation (Betensky, 2015; Li et al., 2009).
Several questions remain to be addressed for the practical implementation of the proposed framework:
  • How should $\theta_{0}$ be defined, taking into account the clinical context and risk tolerance of different stakeholders? For example, patients or clinicians might tolerate greater uncertainty in OS data if a robust clinical benefit is ascertained on an intermediate/surrogate endpoint.
  • How should $\theta^{*}$ be defined given potential confounding by intercurrent events? If a substantial number of patients switch to subsequent therapy after tumor progression, then it may be desirable to relax $\theta^{*}$ since subsequent therapies can lead to a smaller OS difference (Korn et al., 2011).
  • When should OS be evaluated for safety? Should it only be evaluated when OS is sufficiently mature (however we choose to define “mature”)? Or should it also be evaluated at key timepoints (e.g., interim and final analyses of the primary endpoint)?
Another question that may come up is regarding non-proportional hazards (NPH). Is the proposed framework applicable under NPH? Our opinion is yes, that at least a good part of the framework can be applied if the overall HR is still considered a meaningful treatment effect summary; for example, analysis timing can still be based on the number of deaths $d$ and $\theta^{*}$ can still be used as the threshold for signaling overall OS detriment. However, if it is known a priori that a NPH effect could exist, then operating characteristics should be evaluated carefully under potential patterns of NPH in order to ensure that whatever plan that’s proposed for OS safety analysis balances risk appropriately over time. Furthermore, supportive analyses such as piecewise exponential and cure rate models can be prospectively planned to help interpret trial results. This is similar to how some sponsors currently handle NPH for efficacy analysis. They conduct event-driven analyses using the logrank test and cox HR, but consider potential NPH when calculating the required number of events. For example, see Checkmate-274 and Keynote-966 (Bajorin et al., 2021; Kelley et al., 2023).
@@
This paper proposed a supportive decision making framework based on a single criteria $\theta^{*}$. It is worth noting that Rodriguez et al. (2024) considered having a ‘grey zone’, which, if a trial were to fall into, would prompt further OS follow-up. Doing so could have the effect of ruling out substantial OS detriment while limiting rejection of treatments with marginal OS benefit. While we did not consider such an extension here for simplicity of comparison and presentation, we do find this idea of practical interest worth pursuing in future discussions and research.